OpenStudy (anonymous):

find the gradient of the straight line that passes through the points (2,-4) and (-2,5) and hence find the equation of the line clearly showing each step of your method

OpenStudy (callisto):

gradient = \(\frac{y_2-y_1}{x_2-x_1}\) Put \((x_1, y_1)\) = (2,-4) and \((x_2, y_2)\) = (-2,5) into the above formula to find the gradient. Next, To find the equation of the line, use this: \[\frac{y-y_1}{x-x_1} = gradient\]

OpenStudy (anonymous):

using the equation that was found in the question to determine wether or not the point (-1,3) liesbon this point

OpenStudy (callisto):

Put (-1, 3) into the equation and see it the 2 sides are equal..

OpenStudy (anonymous):

ok thanks